# Crowd-SAM: SAM as a Smart Annotator for Object Detection in Crowded Scenes

### v1

@@ Abstract
In computer vision, object detection is an important task that finds its application in many scenarios. However, obtaining extensive labels can be challenging, especially in crowded scenes. Recently, the Segment Anything Model (SAM) has been proposed as a powerful zero-shot segmenter, offering a novel approach to instance segmentation tasks. However, the accuracy and efficiency of SAM and its variants are often compromised when handling objects in crowded and occluded scenes. In this paper, we introduce Crowd-SAM, a SAM-based framework designed to enhance SAM’s performance in crowded and occluded scenes with the cost of few learnable parameters and minimal labeled images. We introduce an efficient prompt sampler (EPS) and a part-whole discrimination network (PWD-Net), enhancing mask selection and accuracy in crowded scenes. Despite its simplicity, Crowd-SAM rivals state-of-the-art (SOTA) fully-supervised object detection methods on several benchmarks including CrowdHuman and CityPersons. Our code is available at https://github.com/FelixCaae/CrowdSAM.
@@ 1 Introduction
Object detection in crowded scenes has emerged as a challenging task in computer vision, widely applied in areas such as traffic management, autonomous driving, and surveillance. The primary focus lies in identifying and locating densely packed common objects like pedestrians, where occlusion and clustering present significant challenges. Significant progress has been made in recent years, including two-stage methods[43, 57] and query-based methods[21, 60, 8]. However, these methods follow a supervised manner and necessitate extensive labeled training samples, incurring a considerable annotation cost of approximately 42.4 seconds per object[36] The density and complexity of crowded scenes further aggravate the annotation burden  [^1] On average, an image contains approximately 22 objects in CrowdHuman [35] and 7 in MS COCO [23].
@@
The high cost of collecting object annotations drives the exploration of alternative approaches such as few-shot learning[42, 37, 32], weakly supervised learning [39, 50], semi-supervised learning [27, 47, 40, 28], and unsupervised learning[24, 6, 46, 1, 20]. The best-performing methods, i.e. semi-supervised object detection (SSOD), leverage the unlabeled data for training and achieve great success on common benchmarks such as PASCAL VOC[14] and COCO[23]. However, SSOD introduces extra complexity such as complicated augmentations and online pseudo-labeling.
@@
In recent times, prompt-based segmentation models have attracted much attention due to their flexibility and scalability. Particularly, the Segment-Anything Model (SAM) [19] can effectively and accurately predict the masks of regions specified by prompts, in the any form of points, boxes, masks, or text descriptions. Recognizing its exceptional potential, researchers have made significant efforts to adapt it for various vision tasks such as medical image recognition[29], remote-sensing[4], industrial defect detection[49], etc.
@@
Despite the huge progress following SAM  [17, 44, 51], most of the current work neglects the potential of applying SAM for object detection tasks in crowded scenes. In this paper, we investigate the potential of applying SAM for detection in crowded scenes with such motivations First, SAM is pre-trained on a very large dataset i.e. SA-1B that contains most of the common objects. It would be reasonable to utilize the learned knowledge to ease the difficulty of labeling massive data and training a brand-new detector. Second, SAM demonstrates robust segmentation ability in handling complicated scenes characterized by occlusion and clustered objects that could be difficult for an object detector trained from scratch.
@@
In this paper, we propose Crowd-SAM, a smart annotator powered by SAM for object detection in crowded scenes. As depicted in  Fig. 1, we introduce a self-promoting approach based on DINOv2 to alleviate the cost of human prompting. Our method employs dense grids equipped with an Efficient Prompt Sampler (EPS) to cover as many objects as possible at a moderate cost. To distinguish the masks from multiple outputs precisely in occluded scenes, we propose a mask selection module, termed Part-Whole Discrimination Network (PWD-Net) that learns to distinguish the output with the highest quality. With a lightweight design and fast training schedule, it achieves considerable performance on benchmarks including CrowdHuman[35] and CityPersons[55].
@@
    We introduce Crowd-SAM - a self-prompted segmentation method for labeling crowded datasets that can produce precise and accurate results with only a few examples.
@@
    We design two novel components of Crowd-SAM, i.e. EPS and PWD-Net which can effectively unleash SAM’s ability on crowded scenes.
@@ 2 Related Work
Object detection. General object detection aims to identify and locate objects and is mainly divided into two categories: one-stage detectors and two-stage detectors. One-stage detectors predict bounding boxes and class scores directly from image features[33, 25, 22], while two-stage detectors first generate region proposals and then classify and refine these proposals [34, 10, 9]. Recently, end-to-end object detectors such as DETR[2, 61, 52] deprecate the hand-crafted modules such as non-maximum suppression (NMS) by adopting one-to-one matching in the training phase, showing great potential in a wide variety of areas.
@@
However, applying these detectors directly to pedestrian detection tasks will lead to performance degradation due to the fact that pedestrians are often crowded and occluded. Early work[30] proposes integrating extra features into a pedestrian detector to explore the low-level visual cues, while follow-up methods [5, 53] attempt to utilize the head areas for better representation learning. In [53], an anchor is associated with two targets, the whole body and the head part, to achieve a more robust detector from joint training. Alternative methods focus on the design of loss functions to improve the training process. For example, RepLoss[43] encourages the prediction consistency of the same target while repels ones from different targets. Recently, Zheng et al. [60] model the relation of queries to improve DETR-based detectors in crowded scenes and achieves remarkable success. Although these works have pushed the boundaries of object detection in crowded scenes to a new stage, they all rely on a large number of labeled samples for training, which is labor-intensive. This limitation inspires us to develop label-efficient detectors and automatic annotation tools, with the help of SAM.
@@
Few-shot object detection (FSOD). The task aims to detect objects of novel classes from limited samples. FSOD methods can be generally categorized into meta-learning based [48, 15] and fine-tuning based methods[42, 37, 32]. Meta-RCNN[48] processes the query and support images in parallel via a siamese network. The query’s region of interest (ROI) features are fused with the class prototypes to effectively transfer knowledge learned from the support set. TFA[42] proposes a simple two-stage fine-tuning method that only fine-tunes the detector’s last layers. FSCE[37] introduces a supervised contrastive loss in the fine-tune stage to mitigate misclassification issues. De-FRCN[32] stops the gradient from the RPN and scales the gradient from the R-CNN, followed by a Prototypical Calibration Block to refine the classification scores.
@@
Segment Anything Models. SAM[19], proposed as a visual foundation model for the segmentation task, is trained on the SA-1B dataset using a semi-supervised learning paradigm. Its exposure to this vast repository of training samples renders it a highly capable class-agnostic model, effectively handling a wide range of objects in the world. Despite its impressive performance in solving segmentation tasks, it suffers from several issues like domain shift, inefficiency, class-agnostic design, etc. HQ-SAM [17] is proposed to improve its segmentation quality by learning a lightweight adapter. Fast-SAM [59] and Mobile-SAM [51] focus on fastening the inference speed of SAM by knowledge distillation. RSprompt[4] enables SAM to generate semantically distinct segmentation results for remote sensing images by generating appropriate prompts. Med-SA[45] proposes a Space-Depth Transpose method to adapt 2D SAM to 3D medical images and a Hyper-Prompting Adapter to achieve prompt-conditioned adaptation. Unfortunately, these approaches necessitate a considerable amount of labeled data for effective adaptation, rendering them impractical for crowded scenes where annotation costs are prohibitive. Different from them, Per-SAM[54] and Matcher [26] teaches SAM to recognize specified objects with only one or few instances by extracting training-free similarity priors. SAPNet [44] proposes a weakly-supervised pipeline for instance segmentation. Although these approaches reduce the data requirement, they still lag behind the demands of crowded scenes, such as pedestrian detection, particularly in occlusion scenes.
@@ 3.1 Preliminaries
SAM [19] is a promising segmentation model that comprises three main components: (a) an image encoder responsible for feature extraction; (b) a prompt encoder designed for encoding the geometric prompts provided by the users; and (c) a lightweight mask decoder that predicts the masks conditioned on the given prompts. Leveraging extensive training data, SAM demonstrates impressive zero-shot segmentation performance across various benchmarks. In particular, SAM utilizes points and boxes as prompts to specify the interested regions.
DINO [3] represents a family of self-supervised vision transformers [7] designed for general-purpose applications. During its training, DINO employs a self-distillation approach akin to BYOL [11], fostering the learning of robust feature representations. DINOv2 [31] enhances the foundation of DINO by integrating additional pre-training tasks, improving its scalability and stability, especially in large models like ViT-H with 1 billion parameters. Thanks to its enhancement, DINOv2 demonstrates strong feature representation ability, especially for the task of semantic segmentation.
3.2 Problem Definition and Framework
As shown in  Fig. 1, our goal is to detect pedestrians in crowded scenes with few annotated data. We formulate this problem as a one-class few-shot detection task. A common few-shot approach is to split the data into the base split and the novel split. Differently, we directly use the data of the target class to train our model, since the foundation models have already been trained on massive data. In particular, we use segmentation masks as intermediate results which can be easily converted to bounding boxes. During the training and evaluation process, only box annotations are provided.
Naive study on SAM auto generator. Before we introduce our method, we take a deeper look into how the prompt number affects SAM’s performance in crowded scenes. To this end, we conduct several naive studies on CrowdHuman[35] with SAM’s auto generator, which utilizes grid points to search every region.  Tab. 1 conveys three key observations: (1) dense grids are necessary for crowded scenes; (2) the ambiguity of point prompts and class-agnostic prompts distribution incur many false positives (FPs); (3) decoding time is a non-negligible burden when grid size is large. It is concluded that the dense prompts and FP removal are key aspects in designing SAM-based methods for detection/segmentation tasks in crowded scenes.
N 16 32 64 128 Oracle
@@ 3.2 Problem Definition and Framework
Table 1: Comparisons of different grid sizes adopted by SAM generator on Crowd-Human[35]. Oracle model derives the prompts by computing the center of the ground truth boxes. T is the decoding time collected on a 3090 Ti GPU card.
Framework. Inspired by the above studies, we equip SAM [19] with several proper components to achieve an accurate and efficient annotation framework, as illustrated in  Fig. 2. To accurately locate pedestrians, we employ the foundational model DINOv2[31] to predict a semantic heat map, a task that can be formulated with a simple binary classifier. To discriminate the output masks that is a mixture of correct masks, backgrounds, and part-level masks, we adopt Part-Whole Discrimination Network that takes as input both the learned tokens from SAM and the semantic-rich tokens from DINOv2 to re-evaluate all the outputs. Finally, to handle the redundancy brought by the use of dense grids, we propose an efficient prompt sampler (EPS) to decode the masks at a moderate cost.
@@
3.3 Class-specific Prompts Generation
Generating a unique point prompt for each pedestrian in the dense scenes is de facto a non-trivial problem. Thus, we take a step back and study how to detect pedestrians with multiple prompts associated with one person, and then use the proper post-processing methods to remove duplicates. To this end, we adopt a heat-map-based prompt generation pipeline that classifies the regions first and generates prompts from the positive regions.
For the input image $I\in\mathbb{R}^{H\times W\times 3}$, we use a pre-trained image encoder $E_{D}$ to extract semantic rich features first. To better transfer the pre-trained features to pedestrian segmentation, we add an MLP block after the final output layer and thus get the adapted features $F_{DINO}\in\mathbb{R}^{{H\over s}\times{W\over s}\times C}$, where $s$ is the patch size of DINOv2 [31] and $C$ is the output channel. Then, we use a segmentation head $\text{Head}_{CLS}$ to classify $F_{DINO}$ pixel by pixel, resulting in a heatmap ^$\hat{\mathbb{H}}\in[0,1]^{{H\over s}\times{W\over s}}$ that indicates the locations of pedestrians as ^$\hat{\mathbb{H}}=\text{Head}_{CLS}(F_{DINO})$.
Given only the bounding box annotations $B\in[0,1]^{N\times 4}$, where $N$ is the number of targets, this binary segmentation head can be optimized with box-level supervision. However, the coarse boundaries would incur considerable points scattering in the background regions. To alleviate this issue, we use SAM [19] to obtain high-quality mask-level pseudo labels, which is illustrated in  Fig. 4, with ground truth (GT). The decoded masks are then merged into a single foreground mask $\mathbb{H}\in\{0,1\}^{256\times 256}$. We use dice loss for training the adapter and segmentation head with the generated pseudo masks as follows:
@@ 3.3 Class-specific Prompts Generation
where $f$ is a up-sampling function that resizes $\mathbb{H}$ to $256\times 256$. During inference, we use a threshold $t$ for mask binarization, which is simply set 0.5 in our experiments. The binarized masks are mapped to point prompts $P_{G}$ which only contain those in positive regions.

### v2

@@ Abstract
Object detection is an important task that finds its application in a wide range of scenarios. Generally, it requires extensive labels for training, which is quite time-consuming, especially in crowded scenes. Recently, Segment Anything Model (SAM) has emerged as a powerful zero-shot segmenter, offering a novel approach to instance segmentation. However, the accuracy and efficiency of SAM and its variants are often compromised when handling objects in crowded scenes where occlusions often appear. In this paper, we propose Crowd-SAM, a SAM-based framework designed to enhance the performance of SAM in crowded scenes with the cost of few learnable parameters and minimal labeled images. We introduce an efficient prompt sampler (EPS) and a part-whole discrimination network (PWD-Net), facilitating mask selection and contributing to an improvement in accuracy in crowded scenes. Despite its simplicity, Crowd-SAM rivals state-of-the-art fully-supervised object detection methods on several benchmarks including CrowdHuman and CityPersons. Our code is available at https://github.com/FelixCaae/CrowdSAM.
@@ 1 Introduction
Object detection in crowded scenes is a fundamental task in areas such as autonomous driving and video surveillance. The primary focus lies in identifying and locating densely packed common objects like pedestrians and vehicles, where occlusions present significant challenges. Great progress has been made in recent years, including two-stage methods [45, 60] and query-based methods [22, 62, 8]. However, these methods mainly follow a supervised manner and necessitate extensive labeled training samples, incurring a considerable annotation cost of approximately 42.4 seconds per object [38] The density and complexity of crowded scenes further aggravate the annotation burden.  [^1] On average, an image contains approximately 22 objects in CrowdHuman [37] and 7 in MS COCO [24]
@@
The high cost of collecting object annotations drives the exploration of alternatives such as few-shot learning [44, 39, 34], weakly supervised learning [41, 53], semi-supervised learning [29, 49, 42, 30], and unsupervised learning [26, 6, 48, 1, 21]. The best-performing ones, i.e. semi-supervised object detection (SSOD), leverage both labeled and unlabeled data for training and achieve a big success on common benchmarks e.g. PASCAL VOC [15] and COCO [24]. Unfortunately, SSOD introduces extra complexity such as complicated augmentations and online pseudo-labeling.
@@
Recently, prompt-based segmentation models have received increasing attention due to their flexibility and scalability. Particularly, Segment Anything Model (SAM) [20] show its high capability to effectively and accurately predict the masks of regions specified by prompts, in any form of points, boxes, masks, or text descriptions. Recognizing its exceptional potential, researchers have made many efforts to adapt it for various vision tasks such as medical image recognition [31], remote sensing analysis [4, 12], industrial defect detection [52], etc.
@@
Despite the huge progress [18, 46, 54] following SAM, applying SAM for object detection in crowded scenes is seldom studied. In this paper, we investigate the potential of SAM in such cases with two motivations. First, SAM is pre-trained on a very large dataset i.e. SA-1B that contains most of the common objects and it is thus reasonable to utilize the knowledge to facilitate labeling massive data and training a brand-new detector. Second, SAM demonstrates a robust segmentation ability in handling complicated scenes characterized by clustered objects that are difficult for an object detector trained from scratch.
@@
To this end, we propose Crowd-SAM, a smart annotator powered by SAM for object detection in crowded scenes. As depicted in Fig. 1, we introduce a self-promoting approach based on DINOv2 to alleviate the cost of human prompting. Our method employs dense grids equipped with an Efficient Prompt Sampler (EPS) to cover as many objects as possible at a moderate cost. To distinguish the masks from multiple outputs precisely in occluded scenes, we design a mask selection module, termed Part-Whole Discrimination Network (PWD-Net) that learns to differentiate the output with the highest quality in Intersection over Union (IoU) score. With a lightweight model design and fast training schedule, it delivers considerable performance on public benchmarks including CrowdHuman [37] and CityPersons [58].
@@
    We propose Crowd-SAM, a self-prompted segmentation method, for labeling images containing clustered objects, producing accurate results with only a few examples.
@@
    We design two novel components of Crowd-SAM, i.e. EPS and PWD-Net, which effectively unleash the ability of SAM on crowded scenes.
@@ 2 Related Work
Object Detection. General object detection aims to identify and locate objects and is mainly divided into two categories: i.e. one-stage detectors and two-stage detectors. One-stage detectors predict bounding boxes and class scores by using image features[35, 27, 23], while two-stage detectors first generate region proposals and then classify and refine them [36, 10, 9]. Recently, end-to-end object detectors e.g. DETR[2, 63, 55] have replaced the hand-crafted modules such as Non-Maximum Suppression (NMS) by adopting one-to-one matching in the training phase, showing great potential in a wide variety of areas.
@@
However, applying these detectors directly to pedestrian detection tends to incur performance degradation due to the fact that pedestrians are often crowded with occlusions appearing. Early work [32] proposes to integrate extra features into a pedestrian detector to explore low-level visual cues, while follow-up methods [5, 56] attempt to utilize the head areas for better representation learning. In [56], an anchor is associated with two targets, the whole body, and the head part, to achieve a more robust detector from joint training. AdaptiveNMS [25] adjusts the NMS threshold by predicting the density of pedestrians. Alternative methods focus on the design of loss functions to improve the training process. For example, RepLoss [45] encourages the prediction consistency of the same target while repels the ones from different targets. Recently, Zheng et al. [62] models the relation of queries to improve DETR-based detectors in crowded scenes and achieves remarkable success. Although these works have pushed the boundaries of object detection in crowded scenes to a new stage, they all rely on a large number of labeled samples for training, which is labor-intensive. This limitation inspires us to develop label-efficient detectors and automatic annotation tools, with the help of SAM.
@@
Few-Shot Object Detection (FSOD). This task aims to detect objects of novel classes with limited training samples. FSOD methods can be roughly classified into meta-learning based [51, 16] and fine-tuning based ones [44, 39, 34]. Meta-RCNN [51] processes the query and support images in parallel via a siamese network. The Region of Interest (RoI) features of the query are fused with the class prototypes to effectively transfer knowledge learned from the support set. TFA [44] proposes a simple two-stage fine-tuning method that only fine-tunes the last layers of the network. FSCE [39] introduces a supervised contrastive loss in the fine-tuning stage to mitigate misclassification issues. De-FRCN [34] stops the gradient from the RPN and scales the gradient from R-CNN[36], followed by a prototypical calibration block to refine the classification scores.
@@
Segment Anything Models. SAM [20], a visual foundation model for segmentation tasks, is trained on the SA-1B dataset using a semi-supervised learning paradigm. Its exposure to this vast repository of training samples renders it a highly capable class-agnostic model, effectively handling a wide range of objects in the world. Despite its impressive performance in solving segmentation tasks, it suffers from several issues like domain shift, inefficiency, class-agnostic design, etc. HQ-SAM [18] is proposed to improve its segmentation quality by learning a lightweight adapter. Fast-SAM [50] and Mobile-SAM [54] focus on fastening the inference speed of SAM by knowledge distillation. RSprompt [4] enables SAM to generate semantically distinct segmentation results for remote sensing images by generating appropriate prompts. Med-SA [47] presents a space-depth transpose method to adapt 2D SAM to 3D medical images and a hyper-prompting adapter to achieve prompt-conditioned adaptation. Unfortunately, these approaches necessitate a considerable amount of labeled data for effective adaptation, making them impractical for crowded scenes where annotation costs are prohibitive. Different from them, Per-SAM [57] and Matcher [28] teach SAM to recognize specified objects with only one or few instances by extracting training-free similarity priors. SAPNet [46] builds a weakly-supervised pipeline for instance segmentation. Although these approaches reduce data requirements, they still lag behind the demands of crowded scenes, such as pedestrian detection, particularly with occlusions.
@@ 3.1 Preliminaries
SAM [20] is a powerful and promising segmentation model that comprises three main components: (a) an image encoder responsible for feature extraction; (b) a prompt encoder designed to encode the geometric prompts provided by users; and (c) a lightweight mask decoder that predicts the masks conditioned on the given prompts. Leveraging extensive training data, SAM demonstrates impressive zero-shot segmentation performance across various benchmarks. In particular, SAM makes use of points and boxes as prompts to specify interested regions.
DINO [3] represents a family of Vision Transformers (ViT) [7] learned in a self-supervised manner designed for general-purpose applications. During its training, DINO employs a self-distillation strategy akin to BYOL [11], fostering the learning of robust representations. DINOv2 [33] strengthens the foundation of DINO by integrating several additional pre-training tasks, improving its scalability and stability, especially for large models e.g. ViT-H (1 billion parameters). Thanks to its enhancement, DINOv2 shows a strong representation ability, in particular for the task of semantic segmentation.
3.2 Problem Definition and Overall Framework
Problem Definition. As shown in Fig. 1, our goal is to detect objects (e.g. pedestrians) in crowded scenes with few annotated data. We formulate this problem as a one-class few-shot detection task. A common few-shot pipeline is to divide data into the base split and the novel split. Differently, we directly use the data of the target class for model training, as the foundation models have already been trained on massive data. In particular, we employ segmentation masks as intermediate results which can be easily converted to bounding boxes. During the training and evaluation processes, only box annotations are provided.
Naive Study on SAM Auto-generator. The prompt number affects the performance of SAM and we analyze this issue for crowded scenes. In this case, we conduct several naive studies on CrowdHuman [37] with the auto generator of SAM, which utilizes grid points to search every region.  Tab. 1 conveys three key observations: (1) dense grids are necessary for crowded scenes; (2) the ambiguity of distributions of point prompts and class-agnostic prompts incur many false positives (FPs); (3) the decoding time is a non-negligible burden when the grid size is large. Therefore, the dense prompts and FP removal are key aspects in designing SAM-based methods for detection/segmentation tasks in crowded scenes.
Table 1: Comparison in terms of recall, average FPs, and decoding time ($T$) of different grid sizes ($N_{G}$) adopted by the SAM generator on CrowdHuman [37]. The oracle model derives the prompts by computing the center of ground truth boxes. The decoding time is collected on a 3090 Ti GPU card.
$N_{G}$ 16 32 64 128 Oracle
@@ 3.2 Problem Definition and Overall Framework
Overall Framework. Inspired by the studies above, we equip SAM [20] with several proper components to achieve an accurate and efficient annotation framework, as illustrated in Fig. 2. To accurately locate clustered objects, we employ the foundational model DINOv2[33] to predict a semantic heat map, a task that can be formulated with a simple binary classifier. To discriminate the output masks that is a mixture of correct masks, backgrounds, and part-level masks, we design Part-Whole Discrimination Network (PWD-Net) that takes as input both the learned tokens from SAM and the semantic-rich tokens from DINOv2 to re-evaluate all the outputs. Finally, to handle the redundancy brought by the use of dense grids, we propose an Efficient Prompt Sampler (EPS) to decode the masks at a moderate cost.
@@
3.3 Class-specific Prompt Generation
Generating a unique point prompt for each object (e.g. pedestrian) in crowded scenes is de facto a non-trivial problem. Thus, we take a step back and study how to detect objects with multiple prompts associated with one object and apply the proper post-processing techniques to remove duplicates. To this end, we adopt a heatmap-based prompt generation pipeline that initially classifies the regions and then generates prompts from the positive regions.
For the input image $I\in\mathbb{R}^{H\times W\times 3}$, we first use a pre-trained image encoder $E_{D}$ to extract semantic rich features. To better transfer the pre-trained features to object segmentation, we add an MLP block after the final output layer and thus obtain the adapted features $F_{DINO}\in\mathbb{R}^{{H\over s}\times{W\over s}\times C}$, where $s$ is the patch size of DINOv2 [33] and $C$ is the output channel. Then, we employ a segmentation head $\text{Head}_{CLS}$ to classify $F_{DINO}$ pixel by pixel, resulting in a heatmap ^$\hat{\mathbb{H}}\in[0,1]^{{H\over s}\times{W\over s}}$ that indicates the locations of objects as ^$\hat{\mathbb{H}}=\text{Head}_{CLS}(F_{DINO})$.
Given only the bounding box annotations $B\in[0,1]^{N_{B}\times 4}$, where $N_{B}$ is the number of targets, this binary segmentation head can be optimized with box-level supervision. However, the coarse boundaries tend to incur considerable points scattering in background regions. To alleviate this issue, we use SAM [20] to generate high-quality mask-level pseudo labels, which is illustrated in Fig. 4, with ground truth (GT). The decoded masks are then merged into a single foreground mask $\mathbb{H}\in\{0,1\}^{256\times 256}$. We use the dice loss for training the adapter and segmentation head with the generated pseudo masks as follows:
@@ 3.3 Class-specific Prompt Generation
where $f$ is an up-sampling function that resizes $\mathbb{H}$ to $256\times 256$. During inference, we add a threshold $t$ for mask binarization, which is simply set at 0.5 in our experiments. The binarized masks are mapped to point prompts $P_{G}$ which only contain those in positive regions.